# HierarchicalPrune: Position-Aware Compression for Large-Scale Diffusion Models

### v1

@@ Abstract
State-of-the-art text-to-image diffusion models (DMs) achieve remarkable quality, yet their massive parameter scale (8-11B) poses significant challenges for inferences on resource-constrained devices. In this paper, we present HierarchicalPrune, a novel compression framework grounded in a key observation: DM blocks exhibit distinct functional hierarchies, where early blocks establish semantic structures while later blocks handle texture refinements. HierarchicalPrune synergistically combines three techniques: (1) Hierarchical Position Pruning, which identifies and removes less essential later blocks based on position hierarchy; (2) Positional Weight Preservation, which systematically protects early model portions that are essential for semantic structural integrity; and (3) Sensitivity-Guided Distillation, which adjusts knowledge-transfer intensity based on our discovery of block-wise sensitivity variations. As a result, our framework brings billion-scale diffusion models into a range more suitable for on-device inference, while preserving the quality of the output images. Specifically, when combined with INT4 weight quantisation, HierarchicalPrune achieves 77.5-80.4% memory footprint reduction (e.g., from 15.8 GB to 3.2 GB) and 27.9-38.0% latency reduction, measured on server and consumer grade GPUs, with the minimum drop of 2.6% in GenEval score and 7% in HPSv2 score compared to the original model. Last but not least, our comprehensive user study with 85 participants demonstrates that HierarchicalPrune maintains perceptual quality comparable to the original model while significantly outperforming prior works.
@@ 1 Introduction
Diffusion-based text-to-image (T2I) synthesis (Song et al. 2021; Stability AI 2023; Lin, Wang, and Yang 2024; Lipman et al. 2023; Karras et al. 2022; Peebles and Xie 2022; Lu et al. 2024) has emerged as a powerful tool for a wide variety of applications, such as the generation of educational content, creative artwork, and UX/UI design prototyping, intensifying demand for deployable billion-parameter diffusion models (DMs). While recent advances, such as Stable Diffusion 3.5 (SD3.5) (Esser et al. 2024) and FLUX (Black Forest Labs 2024), significantly outperform previous generations (SDXL (Podell et al. 2023), SD1.5 (Rombach et al. 2022), and DALLE-2) in image quality and text alignment, they also come with excessive model sizes (8-11B parameters) and compute demands, limiting the accessibility of such advanced models.
At the same time, despite quantitative metrics (e.g., GenEval (Ghosh, Hajishirzi, and Schmidt 2023) and DPG-Bench (Hu et al. 2024b)) suggesting smaller models (approx. 2B parameters) like SANA-Sprint (Chen et al. 2025b) perform better than their large-scale counterparts, user-based evaluations on the Artificial Analysis Leaderboard [^1] https://artificialanalysis.ai/text-to-image/arena?tab=leaderboard-text reveal a significant gap in the perceived quality between compact models and their larger counterparts, which quantitative metrics fail to capture. As such, there is a longstanding need for deploying large models even in resource-constrained settings in order to provide users with high-quality T2I capabilities.
Figure 1: HierarchicalPrune achieves 79.5% memory reduction (left) while maintaining image quality. User study (right) demonstrates minimal quality drop (4.8-5.3%), contrary to the excessive degradation (11.1-52.2%) of prior methods.
@@
Figure 2: High-resolution image samples generated by our pruned/distilled model using HierarchicalPrune, showcasing its superior visual quality across various visual styles, precisely following text prompts, and preserving the ability to draw typography.
Nonetheless, this endeavour constitutes a very challenging task due to the high memory and compute intensity (Li et al. 2025) of large DMs, thereby often requiring cloud solutions equipped with high-end GPUs and a minimum of 80GB VRAM. Concurrently, most of the renowned recent releases of foundational DMs, including SD3.5 (Esser et al. 2024), FLUX (Black Forest Labs 2024), and Seedream 3.0 (ByteDance 2025), are built upon multi-modal diffusion Transformer (MMDiT) backbones (Esser et al. 2024) instead of U-Net, revealing new compression opportunities (Section 2.1). At the moment, existing efforts to improve efficiency face important limitations. Firstly, sampling step reduction (Li et al. 2023; Stability AI 2023; Sauer et al. 2023) and efficient operator design (Xie et al. 2025; Dao et al. 2022) are tailored to improving the speed of DMs rather than reducing memory requirements, leaving the important task of DM deployment in memory-constrained devices unresolved. Secondly, existing depth-pruning methods (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024) show promising results in both memory and computation reduction (Kim et al. 2024a; Fang et al. 2024), outperforming width pruning (Fang, Ma, and Wang 2023; Castells et al. 2024a), but face critical scalability challenges. While they achieve reasonable compression on U-Net-based small DMs (2.6B or less), they fail to compress large-scale, state-of-the-art (SOTA) DMs such as SD3.5 Large (8B) and FLUX (11B) (experiencing significant degradation at 20-30% memory reduction as demonstrated in Table 1). The full-block-removal methods employed therein cannot capture the fine-grained impact of the subcomponents within each block, and more importantly, the different roles of blocks in different positions across the network’s hierarchy, leading to excessive performance drop at high pruning ratios. Lastly, orthogonal to pruning, reduced-precision computation (He et al. 2023; Li et al. 2025; Wang et al. 2024) has been proven effective, but it has not been combined with prior block-removal DM compression methods.
@@
  • Through an extensive user study with 85 participants, we demonstrate that HierarchicalPrune significantly outperforms all the baselines: KOALA and BK-SDM shows a substantial 44.0-52.2% user-perceived quality degradation, a SOTA small-scale DM (SANA-Sprint-1.6B) show a 11.1-14.2% drop. In contrast HierarchicalPrune shows a mere 4.8-5.3% degradation (Fig. 1).
@@ 2.1 Motivation
Examining the architecture and performance of MMDiT-based DMs, we identify a two-fold hierarchy: i) inter-block and ii) intra-block hierarchy. Inter-block hierarchy refers to the organisation of blocks along the overall architecture and implies a functional hierarchy where blocks have different responsibilities based on their position. Intra-block hierarchy focuses on the fact that individual blocks are internally composed of subcomponents, which in the case of MMDiT consist of Norm, Context Norm, Attention, MLP, and Context MLP modules (Esser et al. 2024), with each subcomponent having a varying impact on performance based on its type and position.
Given this dual hierarchy, we conjectured that each block is responsible for different aspects of the generated images and that its position in the DM architecture largely determines these aspects. This, in turn, holds for subcomponents but is also affected by their type. Concretely, we hypothesised that the contribution of different blocks to the output image is not uniform, both in terms of importance to overall performance and influence on specific image traits (e.g., structure, texture or finer details).
To investigate this, we first conducted a contribution analysis on SD3.5 Large Turbo over the HPSv2 dataset (Wu et al. 2023b), by removing both individual MMDiT blocks (Fig. 4a) and subcomponents (Fig. 4b and 4c in this section, and Fig. 8b8e in Appendix B), and comparing the performance before and after. We observe that each of the subcomponents, as well as whole-block removal, demonstrates different patterns of impact at different locations. This was further highlighted when analysing the joint removal of multiple subcomponent types, where different subcomponent combinations exhibited significantly different effects on the final performance (see Appendix B).
@@
Despite this variability, existing full-block removal methods (Lee et al. 2024; Kim et al. 2024a) treat MMDiT blocks homogeneously, compromising performance when targeting high compression ratios. Our observations reveal a critical shortcoming: by ignoring the inter-block differences, these methods inadvertently prune blocks that are disproportionately important to visual quality while retaining less impactful ones. Moreover, by coarsely removing whole blocks, existing methods not only discard redundant layers but also eliminate subcomponents that might be essential for capturing fine-grain features. As shown in Section 3.2, this lack of differentiation leads to a steep drop in model performance under aggressive pruning rates. These insights motivate us to design hierarchy-informed techniques for the effective compression of large MMDiT-based DMs.
@@
Figure 4: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block (a), following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024), or an intra-block subcomponent (b, c and see Fig. 8 in Appendix B for full set of analysis). We report the performance drop compared to the original model. The discrepancy in performance drop patterns reveals the different patterns of importance of each subcomponent.
Figure 5: Impact of removing MMDiT blocks at different positions. Compared to original output (f), removing earlier layers leads to high impact on image structure (a-c), whereas removing later blocks affects mainly fine details (d, e).
@@ 8 $\triangleright$ Given the prunability per block, remove blocks until compression ratio
10
11if $r\geq r_{\textup{thres}}$ then
@@ 16 $\triangleright$ If moderate pruning, use normal distillation
17
18/* - - - Stage 3 - Post-Training Quantisation - - - */
@@ 20 $\triangleright$ Apply PTQ for further compression
Motivated by our findings in Section 2.1, we propose HierarchicalPrune (Fig. 3), a cohesive, multi-stage pruning and distillation approach that respects the hierarchical nature of DMs. HierarchicalPrune introduces three key techniques: i) Hierarchical Position Pruning (HPP), ii) Positional Weight Preservation (PWP), and iii) Sensitivity-Guided Distillation (SGDistill). These techniques operate complementarily with the objective to maximise the model compression rate while maintaining output quality throughout the T2I process. Algorithm 1 presents the complete methodology. Given a pre-trained model $m$, the first stage (lines 1-3) is responsible for producing a pruned model $m_{\text{pruned}}$ with a compression ratio $r$ using HPP (lines 1-3).
In the second stage (lines 4-7), HierarchicalPrune first prepares the pruned model for distillation, by preserving its most sensitive blocks by means of PWP (line 4), yielding the selectively frozen model $m_{\text{pruned,frz}}$. Then, distillation is applied in order to improve the attainable image quality. Specifically, if the target compression ratio $r$ indicates aggressive compression by surpassing a threshold $r_{\text{thres}}$ (line 5), HierarchicalPrune applies SGDistill (line 6), our proposed sensitivity-guided distillation method. In the case of moderate compression, a simpler distillation process is followed (line 8). As a final step, the resulting model is optionally quantised with 4-bit weights and 16-bit activations (line 9), to achieve further compression gains.
@@
Our analysis reveals that in aggressive pruning settings, attempting to update these highly important blocks often proves detrimental to model performance (showing 31.9% average quality reduction even with PWP, see Table 3). Consequently, SGDistill applies inverse distillation weights—assigning minimal or zero update weights to the most important blocks, while concentrating updates on less sensitive components. This approach, combined with subcomponent (e.g., Normalisation and MLP within an MMDiT block) pruning for the remaining portion of the DM after pruning up to $r_{\text{thres}}$ (line 5) with HPP and PWP, extends our hierarchical approach from the block-level to the intra-block subcomponents, preserving the carefully tuned parameters of critical blocks while allowing adaptation in less sensitive regions. Our experimental results confirm this hypothesis, showing that protecting sensitive blocks from significant updates during distillation is key to maintaining quality at high compression ratios, reducing average quality degradation from 31.9% to 10.1% (Table 3).
@@ 3.1 Experimental Setup
Architectures, Datasets, and Implementation: We target SD3.5 Large Turbo (8B) and FLUX.1-Schnell (12B), two SOTA models designed to perform diffusion tasks within a small number of steps (e.g., 4). Following prior work (Lee et al. 2024), we use the YE-POP dataset (HuggingFace 2024) consisting of 500K images, derived from LAION-5B (Schuhmann et al. 2022). We implemented the HierarchicalPrune pipeline in PyTorch, building upon SD3.5 Large Turbo and FLUX.1-Schnell model checkpoints and pipelines from Diffusers (von Platen et al. 2022). For 4-bit weight quantisation, we adopt bitsandbytes (Dettmers et al. 2022, 2023). Note that other quantisation methods, such as SVDQuant (Li et al. 2025), can also be combined with HierarchicalPrune, and we leave such exploration as future work. We aim to release our codebase upon acceptance of the paper.
@@
Baselines: We compare HierarchicalPrune with two prior works related to depth pruning and distillation: i) KOALA (Lee et al. 2024), and ii) BK-SDM (Kim et al. 2024a). BK-SDM proposed block pruning of U-Net-based models using the CLIP score (Hessel et al. 2021), followed by distillation of the pruned model using knowledge from the original model. KOALA follows a similar method, replacing CLIP-score-based importance ranking with scores derived from each block’s input-output cosine similarity. Moreover, we employ the SOTA small-scale DM, SANA (Chen et al. 2025b) as a baseline in our evaluation. [^2] We would like to emphasise that SANA (Chen et al. 2025b) is a specialised on-device model requiring training from scratch, instead of a compression method, such as KOALA, BK-SDM, and HierarchicalPrune.
Metrics: To quantitatively evaluate all methods, we employ the two most recent and representative image quality metrics, GenEval (Ghosh, Hajishirzi, and Schmidt 2023) (used in (Xie et al. 2025; Esser et al. 2024)) and HPSv2 (Wu et al. 2023a) (used in KOALA (Lee et al. 2024)). Furthermore, we conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics. Similar to (Sheynin et al. 2023) and (Dai et al. 2023), we evaluate two aspects: (i) Text Alignment: How well a generated image follows the description of the text prompt, (ii) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors. On the system resource side, we report the measured peak memory usage and latency of running the target models on diverse GPUs, including A6000 (Table 2), GTX 3090 and A100 GPUs (see Appendix Table 5). [^3] Computational overhead analysis of the distillation process as well as contribution analysis is provided in Appendix C.

### v2

@@ Abstract
State-of-the-art text-to-image diffusion models (DMs) achieve remarkable quality, yet their massive parameter scale (8-11B) poses significant challenges for inferences on resource-constrained devices. In this paper, we present HierarchicalPrune, a novel compression framework grounded in a key observation: DM blocks exhibit distinct functional hierarchies, where early blocks establish semantic structures while later blocks handle texture refinements. HierarchicalPrune synergistically combines three techniques: (1) Hierarchical Position Pruning, which identifies and removes less essential later blocks based on position hierarchy; (2) Positional Weight Preservation, which systematically protects early model portions that are essential for semantic structural integrity; and (3) Sensitivity-Guided Distillation, which adjusts knowledge-transfer intensity based on our discovery of block-wise sensitivity variations. As a result, our framework brings billion-scale diffusion models into a range more suitable for on-device inference, while preserving the quality of the output images. Specifically, combined with INT4 weight quantisation, HierarchicalPrune achieves 77.5-80.4% memory footprint reduction (e.g., from 15.8 GB to 3.2 GB) and 27.9-38.0% latency reduction, measured on server and consumer grade GPUs, with the minimum drop of 2.6% in GenEval score and 7% in HPSv2 score compared to the original model. Finally, our comprehensive user study with 85 participants demonstrates that HierarchicalPrune maintains perceptual quality comparable to the original model while significantly outperforming prior works.
@@ 1 Introduction
Diffusion-based text-to-image (T2I) synthesis (2021ICLRPxTIG12RRHS_SDE; stability_ai_sdxl_turbo; 2022arXiv221002747L_FlowMatching; 2022arXiv220600364K_karras; 2022arXiv221209748P_dit; 2024arXiv240212376L_fit) has emerged as a powerful tool for a wide variety of applications, such as the generation of educational content, creative artwork, and UX/UI design prototyping, intensifying demand for deployable billion-parameter diffusion models (DMs). While recent advances, such as Stable Diffusion 3.5 (SD3.5) (esser2024mmdit) and FLUX (blackforestlabs_flux1), significantly outperform previous generations (SDXL (2023arXiv230701952P_sdxl), SD1.5 (2022CVPRLDM), and DALLE-2) in image quality and text alignment, they also come with excessive model sizes (8-11B parameters) and compute demands, limiting the accessibility of such advanced models.
At the same time, despite quantitative metrics, e.g., GenEval (ghosh2023geneval) and DPG-Bench (hu2024ellaequipdiffusionmodels), suggesting smaller models with approx. 2B parameters like SANA-Sprint (chen2025sanasprintonestepdiffusioncontinuoustime) perform better than their large-scale counterparts, user-based evaluations on the Artificial Analysis Leaderboard [^1] https://artificialanalysis.ai/image/leaderboard/text-to-image reveal a significant gap in the perceived quality between compact models and their larger counterparts, which quantitative metrics fail to capture. As such, there is a longstanding need for deploying large models even in resource-constrained settings in order to provide users with high-quality T2I capabilities.
Figure 1: HierarchicalPrune achieves 79.5% memory reduction (left) while maintaining image quality. User study with 85 participants (right) demonstrates minimal quality drop (4.8-5.3%) with 95% confidence intervals, contrary to the excessive degradation (11.1-52.2%) of prior methods.
@@
Figure 2: High-resolution image samples generated by compressed model using HierarchicalPrune, showcasing its superior visual quality across various visual styles, precisely following text prompts, and preserving the ability to draw typography.
Nonetheless, this endeavour constitutes a very challenging task due to the high memory and compute intensity (li2025svdquant) of large DMs, thereby often requiring cloud solutions equipped with high-end GPUs and a minimum of 80GB VRAM. Concurrently, most of the renowned recent releases of foundational DMs, including SD3.5 (esser2024mmdit), FLUX (blackforestlabs_flux1), and Seedream 3.0 (2025seedream3), are built upon multi-modal diffusion Transformer (MMDiT) backbones (esser2024mmdit) instead of U-Net, revealing new compression opportunities (Section 2.1). At the moment, existing efforts to improve efficiency face important limitations. Firstly, sampling step reduction (2023arXiv230600980L_snapfusion; stability_ai_sdxl_turbo; 2023arXiv231117042S_add) and efficient operator design (xie2025sana; FlashAttention) are tailored to improving the speed of DMs rather than reducing memory requirements, leaving the important task of DM deployment in memory-constrained devices unresolved. Secondly, existing depth-pruning methods (lee2024koala; kim2024bksdm; fang2024tinyfusion) show promising results in both memory and computation reduction (kim2024bksdm; fang2024tinyfusion), outperforming width pruning (fang2023structural; Castells_2024_CVPR_ldpruner), but face critical scalability challenges. While they achieve reasonable compression on U-Net-based small DMs (2.6B or less), they fail to compress large-scale, state-of-the-art (SOTA) DMs such as SD3.5 Large (8B) and FLUX (11B) without experiencing significant degradation at 20-30% memory reduction as demonstrated in Table 1. The full-block-removal methods employed therein cannot capture the fine-grained impact of the subcomponents within each block, and more importantly, the different roles of blocks in different positions across the network’s hierarchy, leading to excessive performance drop at high pruning ratios. Lastly, orthogonal to pruning, reduced-precision computation (he2023ptqd; li2025svdquant; 2024arXiv240203666W_quest) has been proven effective, but it has not been combined with prior block-removal DM compression methods.
@@
  • An extensive user study with 85 participants demonstrates that HierarchicalPrune significantly outperforms all the baselines (Fig. 1): KOALA and BK-SDM suffer from a substantial 44.0-52.2% user-perceived quality degradation, and the SOTA small-scale DM, namely SANA-Sprint-1.6B, shows a 11.1-14.2% drop. In contrast, HierarchicalPrune observes a mere 4.8-5.3% degradation.
@@ 2.1 Motivation
Examining MMDiT-based DMs, we identify a two-fold hierarchy: i) inter-block and ii) intra-block hierarchy. Inter-block hierarchy refers to the organisation of blocks along the overall architecture and implies a functional hierarchy where blocks have different responsibilities based on their position. Intra-block hierarchy focuses on the fact that individual blocks are internally composed of subcomponents, which in the case of MMDiT consist of Norm, Context Norm, Attention, MLP, and Context MLP modules (esser2024mmdit), with each subcomponent having a varying impact on performance based on its type and position.
Given this dual hierarchy, we conjectured that each block is responsible for different aspects of the generated images and that its position in the DM architecture largely determines these aspects. This also holds for subcomponents and is affected by their type. Concretely, we hypothesised that the contribution of different blocks to the output image is not uniform, regarding importance to performance and influence on specific image traits (e.g., structure, texture or finer details).
To investigate this, we conducted a contribution analysis on SD3.5 Large Turbo over the HPSv2 dataset (Wu_2023_HPS), by removing both individual MMDiT blocks (Fig. 4a) and subcomponents (Fig. 4b4c and Fig. 8b8e in Appendix B), and comparing the performance before and after. We observe that each of the subcomponents, as well as whole-block removal, demonstrates different patterns of impact at different locations. This was further highlighted when analysing the joint removal of multiple subcomponent types, where different subcomponent combinations exhibited significantly different effects on the final performance (see Appendix B).
@@
Despite this variability, existing full-block removal methods (lee2024koala; kim2024bksdm) treat MMDiT blocks homogeneously, compromising performance when targeting high compression ratios. Our observations reveal a critical shortcoming: by ignoring the inter-block differences, these methods inadvertently prune blocks that are disproportionately important to visual quality while retaining less impactful ones. Moreover, by coarsely removing whole blocks, existing methods not only discard redundant layers but also eliminate subcomponents that might be essential for capturing fine-grain features. As shown in Section 3.2, this lack of differentiation leads to a steep drop in model performance under aggressive pruning rates. These insights motivate us to design hierarchy-informed techniques for the effective compression of large MMDiT-based DMs.
@@
Figure 4: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block (a), following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion), or an intra-block subcomponent (b, c and see Fig. 8 in Appendix B for full set of analysis). We report the performance drop compared to the original model. The discrepancy in performance drop patterns reveals the different patterns of importance of each subcomponent.
Figure 5: Impact of removing MMDiT blocks at different positions. Compared to original outputs (f), removing earlier layers leads to high impact on image structure (a-c), whereas removing later blocks affects mainly fine details (d, e).
@@ 6 $\triangleright$ Given the prunability per block, remove blocks until compression ratio
8if $r\geq r_{\textup{thres}}$ then
@@ 13 $\triangleright$ If moderate pruning, use normal distillation
14
15/* - - - Stage 3 - Post-Training Quantisation - - - */
@@ 17 $\triangleright$ Apply PTQ for further compression
Motivated by our findings in Section 2.1, we propose HierarchicalPrune (Fig. 3), a cohesive, multi-stage pruning and distillation approach that respects the hierarchical nature of DMs. HierarchicalPrune introduces three key techniques: i) Hierarchical Position Pruning (HPP), ii) Positional Weight Preservation (PWP), and iii) Sensitivity-Guided Distillation (SGDistill). These techniques operate complementarily with the objective to maximise the model compression rate while maintaining output quality throughout the T2I process. Algorithm 1 presents the complete methodology. Given a pre-trained model $m$, the first stage (line 1) is responsible for producing a pruned model $m_{\text{pruned}}$ with a compression ratio $r$ using HPP.
In the second stage (lines 2-5), HierarchicalPrune first prepares the pruned model for distillation, by preserving its most sensitive blocks by means of PWP (line 2), yielding the selectively frozen model $m_{\text{pruned,frz}}$. Then, distillation is applied in order to improve the attainable image quality. Specifically, if the target compression ratio $r$ indicates aggressive compression by surpassing a threshold $r_{\text{thres}}$, HierarchicalPrune applies SGDistill (line 3), our proposed sensitivity-guided distillation method. For moderate compression, a simpler distillation process is followed (line 5). As a final step, the resulting model is optionally quantised with 4-bit weights and 16-bit activations (line 6) to obtain additional compression.
@@
Our analysis reveals that in aggressive pruning settings, attempting to update these highly important blocks often proves detrimental to model performance (showing 31.9% average quality reduction even with PWP, see Table 3). Consequently, SGDistill applies inverse distillation weights—assigning minimal or zero update weights to the most important blocks, while concentrating updates on less sensitive components. This approach, combined with subcomponent (e.g., Normalisation and MLP within an MMDiT block) pruning for the remaining portion of the DM after pruning up to $r_{\text{thres}}$ (line 2) with HPP and PWP, extends our hierarchical approach from the block-level to the intra-block subcomponents, preserving the carefully tuned parameters of critical blocks while allowing adaptation in less sensitive regions. Our experimental results confirm this hypothesis, showing that protecting sensitive blocks from significant updates during distillation is key to maintaining quality at high compression ratios, reducing average quality degradation from 31.9% to 10.1% (Table 3).
@@ 3.1 Experimental Setup
Architectures, Datasets, and Implementation: We target SD3.5 Large Turbo (8B) and FLUX.1-Schnell (12B), two SOTA models designed to perform diffusion tasks within a small number of steps (e.g., 4). Following prior work (lee2024koala), we use the YE-POP dataset (yepop) consisting of 500K images, derived from LAION-5B (2022NeurIPSLAION5B). We implemented the HierarchicalPrune pipeline in PyTorch, building upon SD3.5 Large Turbo and FLUX.1-Schnell model checkpoints and pipelines from Diffusers (von-platen-etal-2022-diffusers). For 4-bit weight quantisation, we adopt bitsandbytes (dettmers2022llmint8; dettmers2023qlora). Note that other quantisation methods, such as SVDQuant (li2025svdquant), can also be combined with HierarchicalPrune, and we leave such exploration as future work. We aim to release our codebase upon acceptance of the paper.
@@
Baselines: We compare HierarchicalPrune with two prior works related to depth pruning and distillation: i) KOALA (lee2024koala), and ii) BK-SDM (kim2024bksdm). BK-SDM proposed block pruning of U-Net-based models using the CLIP score (hessel-etal-2021-clipscore), followed by distilling the pruned model using knowledge from the original model. KOALA follows a similar method, replacing CLIP-score-based importance ranking with scores from each block’s input-output cosine similarity. Also, we compare against SANA-Sprint-1.6B (chen2025sanasprintonestepdiffusioncontinuoustime), a SOTA compact model optimised for efficient on-device deployment.
Metrics: To quantitatively evaluate all methods, we employ the two most recent and representative image quality metrics, GenEval (ghosh2023geneval) (used in (xie2025sana; esser2024mmdit)) and HPSv2 (wu2023human) (used in KOALA (lee2024koala)). Furthermore, we conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics. Similar to (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration), we evaluate two aspects: (i) Text Alignment: How well a generated image follows the description of the text prompt, (ii) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors. On the system resource side, we report the measured peak memory usage and latency of running the target models on diverse GPUs, including A6000 (Table 2), GTX 3090 and A100 GPUs (see Appendix Table 5).